Use a model to determine the amount owed on a loan of $5,000 with simple interest of 4.25%. The model was A=5000+5000(0.0425t), where A is the amount (total principal plus interest) required to repay a loan, and t is the time, in years.
What is the slope of this model? Include units in your answer.
How long would it take for the loan amount to reach $7,000? Round to the nearest tenth.

Answer by mananth(16946)

Use a model to determine the amount owed on a loan of $5,000 with simple interest of 4.25%. The model was A=5000+5000(0.0425t), where A is the amount (total principal plus interest) required to repay a loan, and t is the time, in years.
What is the slope of this model? Include units in your answer.
How long would it take for the loan amount to reach $7,000? Round to the nearest tenth.
A=5000+5000(0.0425t)
A= 212.5t +5000
compare with y=mx+c
Slope = m= 212.5
A=7000
A= 212.5t +5000
7000 = 212.50t+5000
2000/212.50=t
t= 9.412 years